OpenStudy (anonymous):

write an equation of a line that passes through the x-intercept 4 and y-intercept -2?? help me please

OpenStudy (anonymous):

x-int = 4 -> point (4, 0) y-int = -2 -> point (0, -2) from those two points can you get the slope? :) @hellokitty_13

OpenStudy (anonymous):

so u just add 0's to them ??

OpenStudy (anonymous):

mmm x intercept means when y equals 0 y intercept means when x equals 0 ... it's referencing the points on the graph

OpenStudy (anonymous):

ohhhhhh thxsss :) :) :) :) :) :)

OpenStudy (anonymous):

u really helped me

OpenStudy (anonymous):

you're welcome ^_^ I'm glad I could help :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

hey can u help on a different 1 plzzzzz

jigglypuff314 (jigglypuff314):

I can :) post your question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

write an equation of the line that has each pair of intercepts. x-intercept: -3 , y- intercept :6

jigglypuff314 (jigglypuff314):

the format for an equation of a line would be like y = mx + b where b = y-intercept and m = slope

OpenStudy (anonymous):

oh but how do u put it together

jigglypuff314 (jigglypuff314):

you would have to find the slope between the x intercept and the y intercept which would be (6-0)/(0-(-3)) = 6/3 = 2 so now you can plug in 6 for b and 2 for m into y = mx + b to get your equation :)

OpenStudy (anonymous):

oh thxx :)

OpenStudy (anonymous):

btw wht grade r in?

jigglypuff314 (jigglypuff314):

me? I'm in 11th grade :)

OpenStudy (anonymous):

daaam im in 8th wht high school do u go to

OpenStudy (anonymous):

??

jigglypuff314 (jigglypuff314):

mmm I go to Immaculata High School in New Jersey

OpenStudy (anonymous):

whatt ur too far away i live in L.A. :(

jigglypuff314 (jigglypuff314):

lol yeah, there's people from all around the world on this site :) I've seen some from india and Pakistan

OpenStudy (anonymous):

ohh im from pakistan

jigglypuff314 (jigglypuff314):

really? cool ^_^

OpenStudy (anonymous):

yea where r u from

OpenStudy (anonymous):

:)

jigglypuff314 (jigglypuff314):

US new jersey :P

OpenStudy (anonymous):

ohh i \

OpenStudy (anonymous):

kool

OpenStudy (anonymous):

i reall y hate math

jigglypuff314 (jigglypuff314):

lol a lot of people do :P but it's really not that hard, and you're pretty good at it! :) just need some more practice

OpenStudy (anonymous):

ohh lol thx

OpenStudy (anonymous):

oh iamma go byee tT.C.